# Prompt Programming: A Platform for Dialogue-based Computational Problem Solving with Generative AI Models

### v1

@@ 1. Introduction
Generative AI has transformed programming and the skills required to excel in this field (Denny et al., 2024c). Students are increasingly turning to AI tools to assist with programming tasks, often without any formal guidance. This highlights a need for educators to actively teach the new skills required to use these tools effectively. Traditionally, introductory courses focused on teaching students to write code by transforming clear problem statements into syntactically and semantically correct code. However, with large language models (LLMs) now capable of solving this step, there is a growing need to emphasize reading, critically evaluating generated code, and crafting precise natural language prompts to achieve desired outcomes.
Several tools have recently been described to help students develop prompt-crafting skills. For instance, one tool expands prompts into high-level pseudocode to help students structure solutions while programming, while another converts prompts into code fragments to generate feedback for code explanation tasks (Kazemitabaar et al., 2024; Smith et al., 2024). A more direct approach is the concept of ‘Prompt Problems’ which challenge students to create single-shot prompts to solve computational tasks presented visually (Denny et al., 2024b). Building on this idea, in the current work we propose several new features that improve the value of this learning-by-prompting approach: dialogue-based interactions, support for problems involving multiple functions, and on-request execution of generated code. These novel features support more authentic interactions with LLMs, promote critical evaluation of code, and enable iterative, real-world problem-solving.
In this paper, we introduce Prompt Programming, a new publicly available platform to teach students how to craft and refine prompts for LLMs. Unlike previous single-shot approaches, it enables multi-turn conversations for iterative refinement and provides on-request execution to test code (Figure 1). The platform also supports multi-function problems, encouraging integrated problem-solving (Figure 2). Through analysis of student interaction data from a large introductory programming course, we evaluate each of these novel features, guided by the following three research questions:
RQ1: How do students’ problem solving approaches for multi-function problems differ from those for single-functions?
@@ 2. Related work
Prompt problems. Denny et al. (Denny et al., 2024b) introduced the idea of ‘Prompt Problems’, a novel exercise where students craft natural language prompts to solve computational tasks presented visually. Accompanied by a prototype tool called Promptly, this approach evaluates code generated from student-written prompts against a test suite. Implemented in CS1 and CS2 courses, their study revealed positive student engagement, with learners appreciating how these exercises foster computational thinking and expose them to new constructs. However, the prototype tool presented in this work supports only single-turn prompts, which limits how well it can be applied to more complex, real-world tasks.
@@ 3.1. Prompt Programming Platform
We developed the Prompt Programming platform to support iterative problem-solving with LLMs, focusing on three core functionalities: multi-function problems, multi-turn conversations, and on-request code execution. These create an environment where students can iteratively refine prompts, critically evaluate model outputs prior to code execution, and develop a systematic approach to solving complex programming problems. By promoting iterative refinement of prompts and deeper engagement with code, the platform fosters critical thinking and problem-solving skills. A web app demonstrating these functionalities is available online, supporting both C and Python. We detail each of the three functionalities below.
@@
Second, the platform provides a chat-like interface for natural, multi-turn conversations, enabling students to iteratively refine their prompts, as shown in Figures 1(b) and 2(b). Each time a student sends a message, the LLM processes it along with the full conversation history to generate context-aware responses. Besides the student’s verbatim message, we prepend a system prompt specifying the programming language and instructing the LLM to exclude testing code or a main function. A ‘Reset’ button allows students to clear the current conversation history and start afresh.
Finally, the Prompt Programming platform includes on-request code execution, enabling students to critically evaluate generated code before running it against hidden test cases. Using the ‘Run Code’ button, students can execute the most recently generated block of code. A counter tracks code execution requests, which can incentivize students to be selective in their use of this feature. This functionality fosters code comprehension and critical judgment by encouraging students to hypothesize code behavior and address potential issues prior to code execution.
@@ 3.2. Study Setup
In order to investigate how our platform supports students in prompt writing and iterative problem solving, we deployed our Prompt Programming platform as part of a large introductory C programming course involving students, taught at Institution anonymized for review. The platform utilized GPT-4o mini (OpenAI, 2024), popular for its accessibility and low-cost usage. The primary goal of the deployment was to collect data on how students used features like multi-turn conversations and on-request code execution to tackle progressively complex problems, including multi-function problems that mirror real-world programming scenarios. To evaluate their engagement and usage patterns, we collected interaction data, including student queries, the LLM’s responses, and code execution requests. Students were also invited to provide optional open-response feedback, providing qualitative insights into their interactions with Prompt Programming.
The study spanned three laboratory assignments (L7, L9, and L10), each lasting one week and each comprising problems created by the authors, who have extensive teaching experience in programming. The problems were chosen to align with the topics students were learning during the course at the time of each laboratory. In L7, students needed to solve all simple single-function problems: counting negatives in an array (see Figure 1), summing even numbers in an array, and finding the index of the last zero in an array. In L9, students needed to solve at least of the more complex single-function problems: sorting a subarray within specified indices, updating a matrix by propagating s across rows and columns, and adding binary numbers represented as arrays. Finally, in L10, students needed to solve at least of the multi-function problems, which required helper functions and structures: password validation (see Figure 2), guiding a robot’s movement on a grid using a structure and functions for turning and advancing, and checking if an expression’s round parentheses are balanced using a stack and corresponding functions. To prevent students from getting sidetracked within a single conversation, we limited conversations to messages for L7 and for L9 and L10 (conversations would need to be reset by the student once reaching these limits).
@@ 4.3. RQ2: Usage of Multi-turn Conversations
Figure 4(a) highlights that students actively leveraged multi-turn conversations across all labs. In L7, most students succeeded within to messages, reflecting the simplicity of the problems. In L9, and more markedly in L10, as problem complexity increased, students made greater use of the multi-turn feature, with more than 25% of the students who solved a problem in L10 requiring more than 5 messages to refine their solutions. This demonstrates that students did not rely on single, monolithic prompts to solve the problems but instead utilized smaller, iterative steps – a more authentic and practical way to interact with generative AI tools.
@@ 4.6. Limitations
Some limitations in our study suggest avenues for future work. First, the testing framework for each multi-function problem currently uses a single driver file that validates all required functions together. This leads to a compilation error if students attempt to run the code before all required functions have been generated through (iterative) prompting. Instead, one could check each component independently, allowing for more modular testing. Second, we did not compare multi-turn and single-turn conversations through an A/B study, which would reveal insights into how each influences success and students’ problem-solving strategies. Finally, we did not analyze how students’ approaches and success correlate with their overall academic performance, leaving this relationship to be explored in future research.
@@ 5. Concluding Remarks
Prompt Programming is publicly available along with a curated set of programming exercises designed for educational use. The platform offers learners an accessible and interactive environment to practice prompt writing and iterative problem-solving in natural language. Future work will focus on expanding the platform’s capabilities by introducing direct code editing features and broadening the variety of exercises to better support diverse learning needs and skill levels. We also aim to explore multi-language interactions and measure the impact of Prompt Programming on student learning outcomes. In the long term, we plan to integrate mechanisms enabling learners to create and share new exercises, fostering a collaborative learning community.
@@ Acknowledgements.
- Doughty et al. (2024) Jacob Doughty et al. 2024. A Comparative Study of AI-Generated (GPT-4) and Human-crafted MCQs in Programming Education. In Proceedings of the Australasian Computing Education Conference (ACE).
@@
- Lee et al. (2023) Unggi Lee et al. 2023. Generative Agent for Teacher Training: Designing Educational Problem-Solving Simulations with Large Language Model-based Agents for Pre-Service Teachers. In NeurIPS’23 Workshop on Generative AI for Education.
@@
- Prather et al. (2023) James Prather et al. 2023. The Robots Are Here: Navigating the Generative AI Revolution in Computing Education. In Proceedings of the Working Group Reports of the Conference on Innovation and Technology in Computer Science Education (ItiCSE).

### v2

@@ 1. Introduction
Generative AI has transformed programming and the skills required to excel in this field (Denny et al., 2024c). Students are increasingly turning to AI tools to assist with programming tasks, often without any formal guidance (Amoozadeh et al., 2024). This highlights a need for educators to actively teach the new skills required to use these tools effectively. Traditionally, introductory courses focused on teaching students to write code by transforming clear problem statements into syntactically and semantically correct code. However, with large language models (LLMs) now capable of solving this step, there is a growing need to emphasize reading, critically evaluating generated code, and crafting precise natural language prompts to achieve desired outcomes.
Several tools were recently described to help students develop prompt-crafting skills. For instance, CodeAid expands prompts into high-level pseudocode to help students structure solutions while programming (Kazemitabaar et al., 2024), whereas an EiPE-inspired tool converts prompts into code fragments to generate feedback for code explanation tasks (Smith et al., 2024). A more direct approach is the concept of ‘Prompt Problems’ which challenge students to create single-shot prompts to solve computational tasks presented visually (Denny et al., 2024b). Building on this idea, in the current work we propose several new features that improve the value of this learning-by-prompting approach: dialogue-based interactions, support for problems involving multiple functions, and on-request execution of generated code. These novel features support more authentic interactions with LLMs, promote critical evaluation of code, and enable iterative, real-world problem-solving.
In this paper, we introduce Prompt Programming, a free web-based platform for students to practice crafting and refining prompts for LLMs, available at https://www.promptprogram.org with support for Python and C-based exercises. Unlike previous single-shot approaches, it enables multi-turn conversations for iterative refinement, provides on-request execution to test code (Figure 1), and supports multi-function problems (Figure 2). We evaluate these novel features using student data from a large introductory programming course, guided by three research questions:
RQ1: What approaches do students employ when solving multi-function problems?
@@ 2. Related work
Prompt problems. Denny et al. (Denny et al., 2024b) introduced the idea of ‘Prompt Problems’, a novel exercise where students craft natural language prompts to solve computational tasks presented visually. Accompanied by a prototype tool called Promptly, this approach evaluates code generated from student-written prompts against a test suite. Implemented in CS1 and CS2 courses, their study revealed positive student engagement, with learners appreciating how these exercises foster computational thinking and expose them to new constructs. However, the prototype tool presented in this prior work supports only single-turn prompts, which limits how well it can be applied to more complex, real-world tasks.
@@ 3.1. Prompt Programming Platform
We developed the Prompt Programming platform to support iterative problem-solving with LLMs, focusing on three core functionalities: multi-function problems, multi-turn conversations, and on-request code execution. These allow students to iteratively refine prompts, critically evaluate model outputs before code execution, and develop a systematic approach to solving complex programming problems. By promoting iterative refinement of prompts and deeper engagement with code, the platform fosters critical thinking and problem-solving skills. We detail each functionality below.
@@
Second, the platform provides a chat-like interface for natural, multi-turn conversations, enabling students to iteratively refine their prompts, as shown in Figures 1(b) and 2(b). Each message is processed with the full conversation history to generate context-aware responses. Besides the student’s verbatim message, we prepend a system prompt specifying the programming language and instructing the LLM to exclude test code or a main function. A ‘Reset’ button lets students clear the conversation history and start afresh.
Finally, the Prompt Programming platform includes on-request code execution, enabling students to critically evaluate generated code before running it against hidden test cases. With the ‘Run Code’ button, students can execute the most recently generated block of code. A counter tracks code execution requests, incentivizing students to be selective in using this feature. This functionality fosters code comprehension and critical judgment by encouraging students to hypothesize code behavior and address issues before running it.
@@ 3.2. Study Setup
In order to investigate how our platform supports students in prompt writing and iterative problem solving, we deployed a version of our Prompt Programming platform, specifically focusing on C exercises, in a large introductory C programming course taught at the University of Auckland with students. The platform used GPT-4o mini (OpenAI, 2024), popular for its accessibility and low cost. The primary goal of the deployment was to collect data on how students used features like multi-turn conversations and on-request code execution to tackle progressively complex problems, including multi-function problems that mirror real-world programming scenarios. To evaluate their engagement and usage patterns, we collected interaction data, including student queries, the LLM’s responses, and code execution requests. Students were also invited to provide optional open-response feedback, providing qualitative insights into their interactions with Prompt Programming.
The study spanned three laboratory assignments (L7, L9, and L10), each lasting one week and each comprising problems created by the authors, who have extensive teaching experience in programming. The problems were chosen to align with the topics students were learning during the course at the time of each laboratory. In L7, students needed to solve all simple single-function problems: counting negatives in an array (see Figure 1), summing even numbers in an array, and finding the index of the last zero in an array. In L9, students needed to solve at least of the more complex single-function problems: sorting a subarray within specified indices, updating a matrix by propagating s across rows and columns, and adding binary numbers represented as arrays. Finally, in L10, students needed to solve at least of the multi-function problems, which required helper functions and structures: password validation (see Figure 2), guiding a robot’s movement on a grid using a structure and functions for turning and advancing, and checking if an expression’s round parentheses are balanced using a stack and corresponding functions. To prevent students from getting sidetracked within a single conversation, we limited conversations to messages for L7, and for L9 and L10 (conversations would need to be reset by the student once reaching these limits).
@@ 4.3. RQ2: Usage of Multi-turn Conversations
Figure 4(a) highlights that students actively leveraged multi-turn conversations across all labs. In L7, most students succeeded within to messages, reflecting the simplicity of the problems. In L9, and more markedly in L10, as problem complexity increased, students made greater use of the multi-turn feature, with more than 25% of the students who solved a problem in L10 requiring more than 5 messages to refine their solutions. This demonstrates that students did not rely on single, monolithic prompts to solve the problems but instead utilized smaller, iterative steps – a more authentic and practical way to interact with generative AI tools (Vaithilingam et al., 2022; Denny et al., 2023).
@@ 4.6. Limitations
Some limitations in our study suggest avenues for future work. First, the testing framework for multi-function problems uses a single driver file that validates all functions together, which can lead to compilation errors if students run code before generating all required components. A more modular approach could enable independent validation of partial solutions. Second, prior exposure to L7 and L9 may have influenced student behavior in L10, such as restart frequency or message refinement. Third, we did not conduct an A/B study comparing single-turn and multi-turn conversations to isolate their impact on problem-solving. Finally, we did not examine how students’ approaches and success relate to academic performance or how reliance on LLM-generated code affects students’ long-term ability to independently generate and understand code. Each of these limitations highlights a direction for future investigation.
@@ 5. Concluding Remarks
Prompt Programming is publicly available along with a curated set of programming exercises designed for educational use. The platform offers learners an accessible, interactive environment to practice prompt writing and iterative problem-solving in natural language. Future work will expand its capabilities by introducing direct code editing and broadening the variety of exercises. We also aim to explore multi-language interactions and measure the impact of Prompt Programming on student learning outcomes. In the longer term, we aim to support educators to customize and assign exercises, and to enable learners to create and share their own problems, fostering a collaborative, community-driven learning environment.
@@ Acknowledgements.
- Amoozadeh et al. (2024) Matin Amoozadeh, Daye Nam, Daniel Prol, Ali Alfageeh, James Prather, Michael Hilton, Sruti Srinivasa Ragavan, and Amin Alipour. 2024. Student-AI Interaction: A Case Study of CS1 Students. In Koli Calling International Conference on Computing Education Research (Koli Calling).
@@
- Doughty et al. (2024) Jacob Doughty, Zipiao Wan, Anishka Bompelli, Jubahed Qayum, Taozhi Wang, Juran Zhang, Yujia Zheng, Aidan Doyle, Pragnya Sridhar, Arav Agarwal, Christopher Bogart, Eric Keylor, Can Kültür, Jaromír Savelka, and Majd Sakr. 2024. A Comparative Study of AI-Generated (GPT-4) and Human-crafted MCQs in Programming Education. In Proceedings of the Australasian Computing Education Conference (ACE).
@@
- Lee et al. (2023) Unggi Lee, Sanghyeok Lee, Junbo Koh, Yeil Jeong, Haewon Jung, Gyuri Byun, Yunseo Lee, Jewoong Moon, Jieun Lim, and Hyeoncheol Kim. 2023. Generative Agent for Teacher Training: Designing Educational Problem-Solving Simulations with Large Language Model-based Agents for Pre-Service Teachers. In NeurIPS’23 Workshop on Generative AI for Education.
@@
- Prather et al. (2023) James Prather, Paul Denny, Juho Leinonen, Brett A. Becker, Ibrahim Albluwi, Michelle Craig, Hieke Keuning, Natalie Kiesler, Tobias Kohn, Andrew Luxton-Reilly, Stephen MacNeil, Andrew Petersen, Raymond Pettit, Brent N. Reeves, and Jaromír Savelka. 2023. The Robots Are Here: Navigating the Generative AI Revolution in Computing Education. In Proceedings of the Working Group Reports of the Conference on Innovation and Technology in Computer Science Education (ItiCSE).